OpenStudy (anonymous):

I need help with this problem Two miner league baseball players got a total of 220 hits. Washington had 8 more hits than sanchez. Find the number of hits for each player

OpenStudy (anonymous):

would i divide 220 from 8

OpenStudy (anonymous):

x+x+8=220 2x+8=220 2x=212 x=106 and then you plug in for x x(Sanchez)=106 x+8(Washington)=114

OpenStudy (anonymous):

Ok I get it now. thanks

OpenStudy (anonymous):

yup no problem